# Death by Impact: Evidence for Merger-Driven Quenching in a Collisional Ring Galaxy at Cosmic Noon

### v1

@@ 1 Introduction
Massive quiescent galaxies observed in the early Universe appear to have undergone rapid cessation of star formation, yet the physical mechanisms responsible for this transition remain a topic of active debate (Man & Belli, 2018, and references therein). A central question is whether all quiescent galaxies experience a starburst phase prior to quenching, and what processes govern such rapid evolutionary changes. The epoch known as Cosmic Noon – characterized by peak levels of star formation, quiescence, and galaxy mergers – provides a crucial window for investigating these dynamics (e.g., Madau & Dickinson, 2014; Conselice, 2014; Förster Schreiber & Wuyts, 2020). Although active galactic nucleus (AGN) feedback is frequently proposed as the primary driver of quenching (e.g., Di Matteo et al., 2005; Somerville et al., 2008; Erb, 2015; Beckmann et al., 2017; Belli et al., 2024; Park et al., 2024; Bugiani et al., 2025), other studies highlight the significant roles of gas depletion and merger-induced perturbations (e.g., Ellison et al., 2024; Suess et al., 2025; Gordon et al., 2025; Hewitt et al., 2025).
@@
One pathway suggests that major, gas-rich mergers act as a central engine for galaxy transformation: they trigger both quasar activity and starbursts, and subsequently quench star formation through feedback-driven gas removal (Hopkins et al., 2008). Alternatively, major mergers may lead to a highly turbulent interstellar medium (ISM) in post-starburst, early-type galaxies, reducing star formation efficiencies by up to two orders of magnitude (e.g., Kretschmer & Teyssier, 2020). Additionally, several studies have suggested that major mergers can suppress star formation efficiency by disrupting the gravitational collapse of gas clouds (e.g., Ellison et al., 2018; van de Voort et al., 2018; Suess et al., 2025). This suppression is often linked to violent disc instability—a dynamical phenomenon that can be triggered in gas-rich major mergers—resulting in elevated ISM turbulence and hindered star formation (e.g., Dekel & Burkert, 2014; Zolotov et al., 2015), even in the absence of a burst. Observational studies have discovered several quenched galaxies involved in major mergers, including in the early universe (e.g., Schreiber et al., 2018; Runco et al., 2022). However, it is often difficult to establish whether the quiescent state of the galaxy is indeed physically linked to the ongoing merger.
@@
One way to address this challenge is to derive a precise timeline of both the merger and the quenching process, and to test whether they are closely associated. This is possible only for special cases. In particular, Collisional Ring Galaxies (hereafter CRG, Lynds & Toomre, 1976; Toomre, 1978; Hernquist & Weil, 1993), which form as a result of mergers, offer precise observational insight into the timeline of the galaxy interaction through detailed analyses of their radially propagating density waves. One of the most prominent examples of CRGs in the local Universe is the Cartwheel Galaxy (Fosbury & Hawarden, 1977; Marcum et al., 1992), widely regarded as the prototypical case of this phenomenon (see also, Marcum et al., 1992; Higdon, 1995; Gao et al., 2003; Wolter & Trinchieri, 2004; Mayya et al., 2005; Murugeshan et al., 2020; Hosseinzadeh et al., 2023). Recent results by Ditrani et al. (2024) show that the star formation rate increases with distance from the center, peaking in the outer ring, while the gas-phase metallicity decreases with radius (Zaragoza-Cardiel et al., 2022). Both trends are consistent with the typical radial behavior observed in disc galaxies (e.g., Guesten & Mezger, 1982; Kennicutt & Evans, 2012; Sánchez et al., 2014; Khoram & Belfiore, 2025). Moreover, Mayya et al. (2023) report that the UV-bright regions within the Cartwheel’s ring host multiple stellar populations, with the bulk of the far-ultraviolet emission arising from nonionizing stars aged between approximately 20 and 150 Myr.
@@
Other studies on CRGs, both in observations and simulations, exhibit enhanced star formation in the ring, typically characterized by a quiescent/low-SFR central region surrounded by a vigorously star-forming ring (e.g., Mayya et al., 2005; Bizyaev et al., 2007; Murugeshan et al., 2020; Ditrani et al., 2024). This trend was specifically observed by Murugeshan et al. (2020) through the analysis of distribution, which in most cases revealed a significant concentration of neutral hydrogen in the ring and a corresponding deficiency in the central regions of CRGs. The observed ring expansion velocities in CRGs vary widely, ranging from a few tens to nearly 250 km/s (e.g., Conn et al., 2011; Parker et al., 2015; Pasha et al., 2025). In previous studies, reconstructing the radial expansion velocity of the ring has enabled estimates of its age and, consequently, the timing of the initial collision—typically within a few hundred million years. Additionally, tidal tails—frequently associated with major mergers in both the nearby and distant Universe (e.g., Wen & Zheng, 2016; Elmegreen et al., 2021)—are present in a subset of these galaxies, offering morphological evidence of recent gravitational interactions (e.g., Conn et al., 2011; Renaud et al., 2018; Mart´ınez-Delgado et al., 2023; Pasha et al., 2025).
@@
Both models and simulations consistently indicate that star-forming rings linked to tidal structures in CRG systems (e.g., Mart´ınez-Delgado et al., 2023) have relatively short lifetimes, typically lasting only 0.2–0.5 Gyr (e.g., Wong et al., 2006; Renaud et al., 2018), and become only faintly detectable, in the local Universe, up to approximately 0.7 Gyr after the initial collision (Wu & Jiang, 2015; Elagali et al., 2018). Moreover, Fiacconi et al. (2012) showed that the ring’s expansion velocity is mainly governed by the mass of the intruding galaxy. Additionally, changes in the SFH can begin as early as 50 Myr prior to the collision, followed by an increase in the Host galaxy’s SFR after the impact (e.g., Renaud et al., 2018).
@@
In this work, we identify a gravitationally bound interacting system at consisting of an AGN-hosting CRG and a massive quiescent companion based on visual inspection of publicly available JWST/NIRCam imaging. By combining high resolution JWST and HST imaging with spatially resolved Keck/MOSFIRE spectroscopy in the Y and H band, we investigate the physical and kinematic properties of the system in detail. This multi-wavelength dataset enables us to derive the SFHs of distinct regions and compare them, for the first time at high redshift, to the timeline of the galaxy collision. Our findings reveal compelling evidence of merger-driven transformation: a starburst episode in the AGN-hosting ring galaxy and simultaneous quenching in the companion. This system presents a rare opportunity to study the impact of both major mergers and AGN activity on galaxy evolution during cosmic noon.
@@
The paper is organized as follows. In Section 2 we introduce the system, followed by a detailed analysis of the spatially resolved photometry in Section 3. In Section 4 we use the spectroscopy to investigate the kinematics and SFHs. We then connect the observed physical transformations in the system to the merger event, and further discuss the implications in Section 5. A summary of the main findings is presented in Section 6. Throughout this paper, we adopt the Planck 2020 cosmology (Planck Collaboration et al., 2020).
@@ 2 The System
We identified an interacting system in the UDS field at redshift 1.61 (see Fig. 1) through visual inspection of publicly available JWST broad-band imaging. The axisymmetric ring structure, along with bright spokes connecting the ring to its central galaxy (UDS 35606, hereafter the Host), are hallmark features of CRGs. The presence of an intruder galaxy (UDS 35616, referred to as the Bullet) at the same redshift, together with a prominent tidal tail, provides strong evidence supporting this classification. Based on dynamical analysis (see Sec. 7), the two galaxies are gravitationally bound and will eventually merge. Spectroscopic confirmation for both the Host and the Bullet is presented in Sec. 4. This system stands out as a rare high-redshift CRG candidate, offering a valuable opportunity for detailed study.
@@
Figure 1 illustrates various morphological features of the system, using JWST/NIRCam imaging at 2.0 and 4.44, as well as an RGB composite. The Bullet is located at a projected distance of kpc from the Host and is visually aligned with the Upper Ring in the background (see Sec. 4.1). Distinct spokes are observed connecting the ring structure to the central Host galaxy. The brightest parts of the ring are more prominent in the bluer bands, while the extended tidal tail is more easily distinguishable in the redder wavelengths. Notably, the Host galaxy is X-ray detected (Chandra observations; Kocevski et al., 2018) and classified as a powerful AGN, with an X-ray luminosity of erg s-1.
For this system, eight JWST/NIRCam broad and medium bands111F090W, F115W, F150W, F200W, F270W, F356W, F410M, F444W. are available from the PRIMER survey (GO 1837; PI: J. Dunlop). Additionally, we utilize mosaics from the 3D-HST survey, covering five HST bands222F606W, F814W, F125W, F140W, F160W. from the ACS and WFC3 instruments (Grogin et al., 2011; Koekemoer et al., 2011; Skelton et al., 2014). The images in each band are resampled and smoothed with a Gaussian kernel accounting for the spatial resolution and the PSF of the reddest band of the same telescope—F160W for HST and F444W for JWST.
We also make use of Keck/MOSFIRE spectra in the Y and H bands from Belli et al. (2017). The observations were conducted with a slit width, yielding a spectral resolution of , with exposure times of 5.5 hours in Y and 1.5 hours in H, in a -seeing. At the redshift of our target, in the rest frame, the Y band optimally captures key absorption features around the Balmer break (–), while the H band (–) contains essential emission lines, including , , and , which are critical for SFR estimation and classification of the ionizing source. The spectra are analyzed in Section 4.
@@ 3 SED Fitting
We identify several regions in the system, shown in the central panel of Fig. 2, using the segmentation map generated by the photutils Python library (Bradley et al., 2024). A region is considered valid if it contains a minimum number of contiguous pixels, each exceeding a defined intensity threshold in the image. This threshold is typically set as 5, i.e. five times the background noise level, to ensure significant detections. The total flux for each photometric band is computed by summing the pixel values within each detected region. The associated uncertainty is derived from two main components: the pixel-wise flux errors obtained from the weight map (combined in quadrature across the region), and the contribution from background flux. To estimate the background, sources are first masked using a 2 detection threshold over the background to exclude faint objects. The median pixel value of the remaining background is then adopted as the nominal background level, and is multiplied by the number of pixels within each region to estimate the total background contamination. Finally, the background contribution and the weight-map-based flux error are added in quadrature. To account for potential systematic uncertainties, a 5-percent systematic error is added to the final uncertainty.
We use prospector (Johnson et al., 2021) to model the SEDs of individual regions, adopting the FSPS stellar population library (Conroy et al., 2009; Conroy & Gunn, 2010), MIST isochrones (Choi et al., 2016), and the C3K spectral library (Cargile et al., 2020), with a Chabrier IMF (Chabrier, 2005). The model incorporates 22 free parameters that describe stars, gas, and dust. For a detailed discussion of the prospector models used here, we refer to Park et al. (2024) and the references therein. The nonparametric SFH is modeled using 14 logarithmically spaced bins, with a Student-T prior applied to (Leja et al., 2019). Two SFH priors are used: the continuity prior (, ) to reduce sharp fluctuations, and the bursty prior (, ) for more flexibility. The two priors yield consistent SFH and SFR estimates, so only results from the continuity prior are presented for clarity.
@@
The Upper Ring exhibits a more complex SFH in comparison to the Lower Ring, which may be partly due to flux contamination from the Bullet and/or the tail, because both the colors and the ionized gas emission lines (analyzed in Sec. 4) suggest that the Upper and Lower Ring have similar SFR. The tidal tail itself is characterized by an old stellar population and signs of recent quenching, with a significantly low SFR (), suggesting a cold-gas poor environment consistent with expectations (e.g., Renaud et al., 2018).
@@ 4.1 Gas Kinematics
In order to analyze the kinematics of the ionized gas, we fit the emission lines in the H-band spectra using the LMFIT package (Newville et al., 2016), performing the fits independently for each region. The continuum is modeled using a second-order polynomial fit to the spectral wavelength window. The emission lines are modeled as Gaussians, all sharing a common velocity and dispersion (), which are fixed to those of the narrow H line. To account for the presence of the AGN and the apparent line broadening in the Host spectrum, we include an additional broad component to the H line. This broad component is allowed to have a central velocity offset of up to km s-1 relative to the narrow component and is constrained to have km s-1. To quantify the uncertainties associated with the line fluxes, we adopt a Monte Carlo simulation technique. This involves perturbing the flux values at each wavelength point 200 times by sampling from a Gaussian distribution, where the mean corresponds to the measured flux and the standard deviation is set by the flux error.
The spectrum of the central region containing the Host galaxy exhibits a prominently broadened H emission line, with a broad component velocity dispersion of km s-1 and a narrow component of km s-1. This, combined with the observed narrow line ratio of [N ii]/H, is characteristic of AGN emission (see, e.g., Kewley et al., 2001; Kauffmann et al., 2003) and is consistent with the classification of the Host galaxy as an X-ray detected AGN (see Sec. 1). The broad component is negligible in both the upper and lower regions.
@@
Having measured the kinematics of the Upper and Lower Ring, we can now attempt to determine the rotational and radial (expansion) velocities of the collisional ring. This can be done by analyzing the line-of-sight (LOS) velocity as it varies with deprojected position angle (PA) at a constant radius along the ring (see Fig. 8). We estimate the ring diameter to be approximately 19.2 kpc (see Appendix B), with an inclination angle of degrees, derived from the observed ring width of 1.6 kpc. The inclination allows for deprojecting the observed LOS velocities. Serendipitously, the lower part of the MOSFIRE slit includes the intersection of the Lower Ring with the system’s major axis (see Appendix B), where the LOS velocity predominantly reflects rotational motion (see Fig. 9). This alignment enables us to derive a deprojected rotational velocity of km/s, fully compatible with the rotational velocity expected for the measured stellar mass (e.g., Cannarozzo et al., 2020). Using the average angular offset of the Upper Ring from the system’s measured PA, estimated to be approximately , we disentangle two perpendicular velocity components—radial (expansion) and rotational—in the Upper Ring (see Fig. 9). Given this geometry, the rotational contribution to the observed H velocity (see Table 2) is estimated to be km s-1. Subtracting this from the measured LOS velocity yields the projected expansion velocity. Accounting for the inclination of the system, we derive a deprojected radial (expansion) velocity of km s-1.

### v2

@@ 1 Introduction
Massive quiescent galaxies observed in the early Universe appear to have undergone rapid cessation of star formation, yet the physical mechanisms responsible for this transition remain a topic of active debate (2018Man, and references therein). A central question is whether all quiescent galaxies experience a starburst phase prior to quenching, and what processes govern such rapid evolutionary changes. The epoch known as Cosmic Noon – characterized by peak levels of star formation, quiescence, and galaxy mergers – provides a crucial window for investigating these dynamics (e.g., 2014Madau; 2014Conselice; 2020AForsterSchreibe). Although active galactic nucleus (AGN) feedback is frequently proposed as the primary driver of quenching (e.g., 2005DiMatteo; 2008MNRASomerville; 2015Erb; 2017Beckmann; 2024Belli; 2024Park; 2025Bugiani), other studies highlight the significant roles of gas depletion and merger-induced perturbations (e.g., 2024Ellison; 2025Suess; 2025Gordon; 2025Hewitt).
@@
Major, gas-rich mergers have been proposed as an important driver of galaxy transformation, potentially operating through multiple pathways. They trigger both quasar activity and starbursts, and subsequently quench star formation through feedback-driven gas removal (2008Hopkins). Alternatively, major mergers may lead to a highly turbulent interstellar medium (ISM) in post-starburst, early-type galaxies, reducing star formation efficiencies by up to two orders of magnitude (e.g., 2020MNRASKretschmer). Additionally, several studies have suggested that major mergers can suppress star formation efficiency by disrupting the gravitational collapse of gas clouds (e.g., 2018Ellison; 2018vandeVoort; 2025Suess). This suppression is often linked to violent disc instability—a dynamical phenomenon that can be triggered in gas-rich major mergers—resulting in elevated ISM turbulence and hindered star formation (e.g., 2014Dekel&Burkert; 2015Zolotov), even in the absence of a starburst. Observational studies have discovered several quenched galaxies involved in major mergers, including in the early universe (e.g., 2018Schreiber; 2022Runco). However, it is often difficult to establish whether the quiescent state of the galaxy is indeed physically linked to the ongoing merger.
@@
The difficulty of establishing a direct causal link between mergers and quenching can be overcome by deriving a precise timeline of both the merger and the quenching process, and then testing whether they are closely associated. This is possible only for special cases. In particular, Collisional Ring Galaxies (hereafter CRG, 1976Lynds; 1978Toomre; 1993Hernquist), which form as a result of mergers, offer precise observational insight into the timeline of the galaxy interaction through detailed analyses of their radially propagating density waves. One of the most prominent examples of CRGs in the local Universe is the Cartwheel Galaxy (1977Fosbury; 1992Marcum), widely regarded as the prototypical case of this phenomenon (see also, 1992Marcum; 1995Higdon; 2003Gao; 2004Wolter; 2005Mayya; 2020Murugeshan; 2023Hosseinzadeh). Recent results by 2024Ditrani show that the star formation rate increases with distance from the center, peaking in the outer ring, while the gas-phase metallicity decreases with radius (2022Zaragoza-Cardiel). Both trends in SFR and metallicity are consistent with the typical radial behavior observed in disc galaxies (e.g., 1982Guesten; 2012Kennicutt; 2014Sanchez; 2025Khoram), confirming that the collisional ring is formed from the gas in the outer disk. Moreover, 2023Mayya report that the UV-bright regions within the Cartwheel’s ring host multiple stellar populations, with the bulk of the far-ultraviolet emission arising from nonionizing stars aged between approximately 20 and 150 Myr. Note that the Cartwheel system is not an isolated case but instead exemplifies a broader class of CRGs in the local Universe.
@@
Other studies on CRGs, both in observations and simulations, exhibit enhanced star formation in the ring, typically characterized by a quiescent/low-SFR central region surrounded by a vigorously star-forming ring (e.g., 2005Mayya; 2007Bizyaev; 2020Murugeshan; 2024Ditrani). This trend was specifically observed by 2020Murugeshan through the analysis of distribution, which in most cases revealed a significant concentration of neutral hydrogen in the ring and a corresponding deficiency in the central regions of CRGs. The observed ring expansion velocities in CRGs vary widely, ranging from a few tens to nearly 250 km/s (e.g., 2011Conn; 2015Parker; 2025Pasha). In previous studies, reconstructing the radial expansion velocity of the ring has enabled estimates of its age and, consequently, the timing of the initial collision—typically within a few hundred million years. Additionally, tidal tails—frequently associated with major mergers in both the nearby and distant Universe (e.g., 2016Wen; 2021Elmegreen)—are present in a subset of these galaxies, offering morphological evidence of recent gravitational interactions (e.g., 2011Conn; 2018Renaud; 2023AMartinez-Delgado; 2025Pasha).
@@
Both models and simulations consistently indicate that star-forming rings linked to tidal structures in CRG systems (e.g., 2023AMartinez-Delgado) have relatively short lifetimes, typically lasting only 0.2–0.5 Gyr (e.g., 2006Wong; 2018Renaud), and become only faintly detectable, in the local Universe, up to approximately 0.7 Gyr after the initial collision (2015Wu; 2018Elagali). Moreover, 2012FiacconiF showed that the ring’s expansion velocity is mainly governed by the mass of the intruding galaxy. Additionally, changes in the SFH can begin as early as 50 Myr prior to the collision, followed by an increase in the Host galaxy’s SFR after the impact (e.g., 2018Renaud).
@@
In this work, we identify a gravitationally bound interacting system at consisting of an AGN-hosting CRG and a massive quiescent companion based on visual inspection of publicly available JWST/NIRCam imaging. By combining high resolution JWST and HST imaging with spatially resolved Keck/MOSFIRE spectroscopy in the Y and H band, we investigate the physical and kinematic properties of the system in detail. This multi-wavelength dataset enables us to derive the SFHs of distinct regions and compare them, for the first time at high redshift, to the timeline of the galaxy collision. Our findings reveal compelling evidence of merger-driven transformation: a starburst episode in the AGN-hosting ring galaxy and simultaneous quenching in the companion. This system presents a rare opportunity to examine how SFHs respond to dramatic galactic perturbations driven by major mergers and AGN activity during cosmic noon.
@@
The paper is organized as follows. In Section 2 we introduce the system, followed by a detailed analysis of the spatially resolved photometry in Section 3. In Section 4 we use the spectroscopy to investigate the kinematics and SFHs. We then connect the observed physical transformations in the system to the merger event, and further discuss the implications in Section 5. A summary of the main findings is presented in Section 6. Throughout this paper, we adopt the Planck 2020 cosmology (2020Planck).
@@ 2 The System
We identified an interacting system in the UDS field at redshift 1.61 (see Fig. 1) through visual inspection of publicly available JWST broad-band imaging. The axisymmetric ring structure, along with bright spokes connecting the ring to its central galaxy (UDS 35606, hereafter the Host), are hallmark features of CRGs. The presence of an intruder galaxy (UDS 35616, referred to as the Bullet) at the same redshift, together with a prominent tidal tail, provides strong evidence supporting this classification. Based on dynamical analysis (see Appendix A), the two galaxies are gravitationally bound and will eventually merge. Spectroscopic confirmation for both the Host and the Bullet is presented in Sec. 4. This system stands out as a rare high-redshift CRG candidate, offering a valuable opportunity for detailed study.
@@
Figure 1 illustrates various morphological features of the system, using JWST/NIRCam imaging at 2.0 and 4.44, as well as an RGB composite. The Bullet is located at a projected distance of kpc from the Host and is visually aligned with the Upper Ring in the background (see Sec. 4.1). Distinct spokes are observed in the F200W-band image, connecting the ring structure to the central Host galaxy. The brightest parts of the ring are more prominent in the bluer bands, while the extended tidal tail is more easily distinguishable in the redder wavelengths. Notably, the Host galaxy is X-ray detected (Chandra observations; 2018Kocevski_UDSXRay) and classified as a powerful AGN, with an X-ray luminosity of erg s-1.
For this system, eight JWST/NIRCam broad and medium bands111F090W, F115W, F150W, F200W, F270W, F356W, F410M, F444W. are available from the PRIMER survey (GO 1837; PI: J. Dunlop). Additionally, we utilize mosaics from the 3D-HST survey, covering five HST bands222F606W, F814W, F125W, F140W, F160W. from the ACS and WFC3 instruments (2011Grogin; 2011Koekemoer; 2014Skelton). The images in each band are resampled and smoothed with a Gaussian kernel accounting for the spatial resolution and the PSF of the reddest band of the same telescope—F160W for HST and F444W for JWST.
We also make use of Keck/MOSFIRE spectra in the Y and H bands from 2017Belli. The observations were conducted with a slit width, yielding a spectral resolution of , with exposure times of 5.5 hours in Y and 1.5 hours in H, in a -seeing. At the redshift of our target, in the rest frame, the Y band optimally captures key absorption features around the Balmer break (–), while the H band (–) contains essential emission lines, including , , and , which are critical for SFR estimation and classification of the ionizing source. The spectra are analyzed in Section 4.
@@ 3 SED Fitting
We identify several regions in the system, shown in the central panel of Fig. 2, using the segmentation map generated by the photutils Python library (larry_bradley_2024_13989456). Given the broad color range of the system (see Fig. 1), the segmentation map is constructed from a stacked combination of all NIRCam images to ensure that all bright regions are included. A region is considered valid if it contains at least ten contiguous pixels, lying above 30 relative to the background noise level, thereby ensuring significant detections. The total flux for each photometric band is computed by summing the pixel values within each detected region. The associated uncertainty is derived from two main components: the pixel-wise flux errors obtained from the weight map (combined in quadrature across the region), and the contribution from background flux. To estimate the background, sources are first masked using a 2 detection threshold over the background to exclude faint objects. The median pixel value of the remaining background is then adopted as the nominal background level, and is multiplied by the number of pixels within each region to estimate the total background contamination. Finally, the background contribution and the weight-map-based flux error are added in quadrature. To account for potential systematic uncertainties, a 5-percent systematic error is added to the final uncertainty.
We use prospector (2021Johnson) to model the SEDs of individual regions, adopting the FSPS stellar population library (2009Conroy; 2010Conroy), MIST isochrones (2016Choi), and the C3K spectral library (2020Cargile), with a Chabrier IMF (2005Chabrier). The model incorporates 22 free parameters that describe stars, gas, and dust. For a detailed discussion of the prospector models used here, we refer to 2024Park and the references therein. The nonparametric SFH is modeled using 14 logarithmically spaced bins, with a Student-T prior applied to (2019Leja). Two SFH priors are used: the continuity prior (, ) to reduce sharp fluctuations, and the bursty prior (, ) for more flexibility. The two priors yield consistent SFH and SFR estimates, so only results from the continuity prior are presented for clarity.
@@
The Upper Ring exhibits a more complex SFH in comparison to the Lower Ring, which may be partly due to flux contamination from the Bullet and/or the tail, because both the colors and the ionized gas emission lines (analyzed in Sec. 4) suggest that the Upper and Lower Ring have similar SFR. The tidal tail itself is characterized by an old stellar population and signs of recent quenching, with a significantly low SFR (), suggesting a cold-gas poor environment consistent with expectations (e.g., 2018Renaud). We note that modest changes to the segmentation boundaries do not introduce significant differences in the derived stellar masses or star formation histories, with results remaining consistent within the quoted uncertainties.
@@ 4.1 Gas Kinematics
In order to analyze the kinematics of the ionized gas, we fit the emission lines in the H-band spectra using the LMFIT package (2016Newville), performing the fits independently for each region. The continuum is modeled using a second-order polynomial fit to the spectral wavelength window. The emission lines are modeled as Gaussians, all sharing a common velocity and dispersion (), which are fixed to those of the narrow H line. To account for the presence of the AGN and the apparent line broadening in the Host spectrum, we include an additional broad component to the H line. This broad component is allowed to have a central velocity offset of up to km s-1 relative to the narrow component and is constrained to have km s-1. To quantify the uncertainties associated with the line fluxes, we adopt a Monte Carlo simulation technique. This involves perturbing the flux values at each wavelength point 200 times by sampling from a Gaussian distribution, where the mean corresponds to the measured flux and the standard deviation is set by the flux error.
The spectrum of the central region containing the Host galaxy exhibits a prominently broadened H emission line, with a broad component velocity dispersion of km s-1 and a narrow component of km s-1. This, combined with the observed narrow line ratio of [N ii]/H, is characteristic of AGN emission (see, e.g., 2001Kewley; 2003Kauffmann) and is consistent with the classification of the Host galaxy as an X-ray detected AGN (see Sec. 1). The broad component is negligible in both the upper and lower regions.
@@
Having measured the kinematics of the Upper and Lower Ring, we can now attempt to determine the rotational and radial (expansion) velocities of the collisional ring. This can be done by analyzing the line-of-sight (LOS) velocity as it varies with deprojected position angle (PA) at a constant radius along the ring (see Fig. 8). We estimate the ring diameter to be approximately 19.2 kpc (see Appendix B), with an inclination angle of degrees, derived from the observed ring width of 1.6 kpc. The inclination allows for deprojecting the observed LOS velocities. Serendipitously, the lower part of the MOSFIRE slit includes the intersection of the Lower Ring with the system’s major axis (see Appendix B), where the LOS velocity predominantly reflects rotational motion (see Fig. 9). This alignment enables us to derive a deprojected rotational velocity of km/s, fully compatible with the rotational velocity expected for the measured stellar mass (e.g., 2020Cannarozzo). Using the average angular offset of the Upper Ring from the system’s measured PA, estimated to be approximately , we disentangle two perpendicular velocity components—radial (expansion) and rotational—in the Upper Ring (see Fig. 9). Given this geometry, the rotational contribution to the observed H velocity (see Table 2) is estimated to be km s-1. Subtracting this from the measured LOS velocity yields the projected expansion velocity. Accounting for the inclination of the system, we derive a deprojected radial (expansion) velocity of km s-1.